# Data Compression for Fast Online Stochastic Optimization

### v1

@@ 8 Numerical examples
  • Online clustering (our memory-efficient method). We use our online clustering algorithm, as described in Algorithm 3, to update cluster assignments. This method discards the data-points once seen, and is thus memory-efficient. Upon constructing the empirical clustered distribution, we solve (8) for the online solution.
  • $k$-means reclustering (our method). Same as above, except using $k$-means clustering at each time step, warm-starting at the previous centers. The procedure is described in Algorithm 2.
  • Wasserstein DRO. At time $t$, we solve (8) with the ambiguity set $\mathcal{P}_{t-1}$.
  • Sample average approximation SAA. At time $t$, we solve the stochastic optimization problem with respect to the empirical measure $\hat{{\mathbf{P}}}_{t-1}$ of the training data. We denote this in-sample value as $H^{\rm SAA}_{t}$.
For all approaches, we use the same training (in-sample) and testing (out-of-sample) datasets. The testing dataset is of size 200, and we consider up to $T=2000$ time steps. For the DRO methods, we select the radii $\varepsilon_{t}$ using the procedure given in Section 7. Specifically, we choose the best $\varepsilon_{t}$ determined through 30 repetitions of the experiment. We compare the following metrics, with values averaged over all repetitions. We additionally show the 25-75th percentiles using shaded regions.
  • In-sample certificate. For all approaches, we compare the in-sample certificates obtained with respect to the training data. For nominal Wasserstein DRO, the certificate is the in-sample objective $H_{t}$. For our online methods, the certificate is the in-sample objective plus the clustering discrepancy, $H^{K}_{t}+\underline{\psi}_{t}^{K}$.
  • Empirical confidence. The empirical confidence is computed as the probability that the in-sample certificate upper bounds the out-of-sample expected value, which is the stochastic optimization objective computed using the empirical measure of the testing dataset, at the given solutions. This value is denoted $1-\hat{\beta}_{t}$.
@@
  • Regret. For our methods, we compute the dynamic regret $R(T,K)$ (9), compared with the theoretical upper bound given in Lemma 4.1.
@@ 8.1 Sparse portfolio optimization
Similarly as in [56, 36], we consider a market that forbids short-selling and has $d$ assets. The uncertain parameters are the daily returns of these assets, given by the random vector $r=(r_{1},\dots,r_{d})\in{\mbox{\bf R}}^{d}$. The percentage weights (of the total capital) invested in the assets are given by the decision vector ${x=(x_{1},\dots,x_{n})\in{\mbox{\bf R}}^{n}}$. We restrict our selection to at most $\gamma$ assets, given by the 00-th norm cardinality constraint below. The distribution ${\mathbf{P}}$ is unknown, but we have access to a streaming dataset $\mathcal{D}_{t}$, updated each day with a new returns vector. Our objective is to minimize the CVaR with respect to variable $x$,
minimize $\begin{array}[]{ll}\text{minimize }&\mathop{\bf CVaR}(-r^{T}x,\alpha)\\ \text{subject to}&\mathbf{1}^{T}x=1,\quad x\geq 0,\quad\|x\|_{0}\leq\gamma,\end{array}$
[…]
@@
We take stock data from the past 5 years of S&P500 (1/1/2020 to 1/1/2025) daily returns, and generate synthetic data from their fitted general Pareto distributions, with correlations preserved using a Gaussian copula. These distributions are chosen to model the data more accurately than a Gaussian fit. We let $\alpha=20\%$, $d=50$ stocks, and restrict our portfolio to at most $\gamma=8$ stocks. We initialize with a dataset of size $n_{0}=5$, and show results for $K=15$ and $K=25$. This range is obtained using the elbow method on the initial dataset, and allowing for adjustments.
@@ Results
In Figure 3, we show the various certificates given in equation (14), and note that the relationships follow the hierarchy presented. The data-compressed in-sample objective value for this maximum-of-affine problem is lower than the in-sample objective of nominal DRO, but adding the clustering discrepancy makes it a valid certificate.
Figure 3: Sparse portfolio, $K=25$. The radius $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$ for all methods. Left: certificates from (14). Right: dynamic regret (9), compared to the theoretical upper bounds.
In Figure 3, we also show the dynamic regret $R(T,K)$ and the regret bound as calculated in Lemma 4.1. As $T\to\infty$, the upper bound is expected to converge to a fixed function of the clustering discrepancy; indeed, we observe that it coincides with $\Phi^{K}_{t}$. This is in line with the theory; since we are solving a maximum-of-affine problem, the term $\bar{\psi}^{K}_{t}$ is reduced to 0. The limiting distance becomes $\Phi^{K}_{t}$, which, as we show in Figure 4 below, is a tighter bound than the Wasserstein-1 distance. This value should converge to its theoretical limit $\Phi_{\star}^{K}$.
Figure 4: Sparse portfolio, $K=25$, $T=10000$, at $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$. Left: in-sample certificates. Right: clustering distances given in Definition 3.4.
In Figure 4, we show the results of our online approaches for a longer horizon, up to $T=10000$, with $\tau=8500$. We observe that the $p=1$ and $p=2$ clustering distances, $d^{K}_{t,1}$ and $(D^{K}_{t,2})^{2}$, as well as $\Phi^{K}_{t}$, are converging, and the out-of-sample performances are also converging in mean. As remarked above, we observe that $d^{K}_{t,1}$ upper bounds $\Phi^{K}_{t}$ without even multiplying by the maximum Lipschitz constant (which is $51/\alpha=$ 5). This illustrates the relative optimality of $\Phi^{K}_{t}$, as opposed to the Wasserstein-1 distance, as proven in Theorem 3.9. In fact, although we have not assumed (or achieved) optimal clustering-induced coupling, the result still held.
@@ 9 Conclusions
We have introduced an online data compression framework for solving Wasserstein DRO problems with streaming data. Our method constructs adaptive ambiguity sets using online clustering, allowing the uncertainty model to evolve as new data arrives, while maintaining out-of-sample performance guarantees. We analyzed the impact of data compression on solution quality, providing finite-sample and asymptotic performance bounds, as well as a sublinear regret analysis with respect to the full-information DRO solution. The framework is compatible with a broad class of clustering algorithms and supports efficient, memory-aware implementations. Numerical experiments in sparse portfolio optimization demonstrate significant computational savings with minimal loss in solution quality.
@@ Acknowledgments
I. Wang and B. Stellato are supported by the NSF CAREER Award ECCS-2239771. I. Wang is also supported by the Princeton Wallace Memorial Fellowship. B. Stellato is also supported by the ONR YIP Award N000142512147. The authors are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University which is a consortium of groups led by the Princeton Institute for Computational Science and Engineering (PICSciE) and Research Computing.
@@ Appendix A Appendices
For all $t$, let $\tilde{\mathbf{Q}}_{t}\in\tilde{\mathbf{B}}_{\varepsilon_{t}}^{p}(\hat{{\mathbf{P}}}_{t})$ and $\mathbf{Q}_{t}\in\mathcal{P}_{t}$. By [40, Lemma 3.7], we observe ${\mathbf{P}}^{\infty}\{\lim_{t\to\infty}\tilde{\mathbf{Q}}_{t}\rightarrow{\mathbf{P}}\}=1$ and ${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\mathbf{Q}_{t}\rightarrow{\mathbf{P}}\right\}=1$. The result follows. ∎
@@
Recall that under the radius $\varepsilon_{t-1}$, the solution $x_{t}$ and optimal value $H_{t}=\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]$ of the DRO problem (4) imply the finite sample performance guarantee (5). We then observe
$\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]$ $\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\Phi^{K}_{t-1},$
with $\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]=H^{K}_{t}$ in the final clause. The first inequality follows from the optimality of $x_{t}$, the second follows from [56, Theorem 5], and $\Phi^{K}_{t}$ is defined in Definition 3.4. From [56, Theorem 4], we note that $\Phi^{K}_{t-1}=0$ for concave functions $f$i.e., when $J=1$. We also observe, using Kantorovich-Rubinstein duality [33],
$\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]$ $\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+W_{1}(\hat{{\mathbf{P}}}_{t-1},\hat{{\mathbf{P}}}^{K}_{t-1}))$
$\leq H^{K}_{t}+\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1}).$
Combining the above relations, we obtain
${\mathbf{P}}^{n_{t-1}}\left(H_{\star}\leq H_{t}\leq\hat{H}^{K}_{t}\leq H^{K}_{t}+\min\left\{\Phi^{K}_{t-1},\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1})\right\}\right)\geq 1-\beta_{t-1}.$
To derive (13), we also note that, with probability $1-\beta_{t-1}$,
$H_{\star}=\min_{x\in\mathcal{X}}\mathbf{E}_{\mathbf{P}}[f(u,x)]\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]\leq H^{K}_{t}+\Phi^{K}_{t-1}.$
Now for the upper bound, from the optimality of $H^{K}_{t}$ and [56, Theorem 5],
$H^{K}_{t}\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]$ $\leq H_{t}+(\tilde{H}_{t}-H_{t})+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2}$
$=H_{t}+\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2},$
where $\Delta_{t}$ is given in Lemma 3.2. By applying Kantorovich-Rubinstein duality,
$H^{K}_{t}\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]\leq H_{t}+\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1}).$
Combining both inequalities, we obtain
$H^{K}_{t}\leq H_{t}+\min\left\{\Delta_{t}+{\max_{j\leq J}(L_{j}/2)}(D^{K}_{t-1,2})^{2},\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+d^{K}_{t-1,1})\right\}.$
Combining the upper and lower bounds gives the desired result.Rearranging terms gives equation (12). ∎
A.2 Proof of the regret bound in Section 4
@@
For brevity, let us denote the following shorthands: $\bar{f}_{t}(x)=\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x)]$, $\bar{f}^{K}_{t}(x)=\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t-1}}\mathbf{E}_{\mathbf{Q}}[f(u,x)],$ and the corresponding solutions $x_{t}=\mathop{\rm argmin}_{x\in\mathcal{X}}\bar{f}_{t}(x)$, $x^{K}_{t}=\mathop{\rm argmin}_{x\in\mathcal{X}}\bar{f}^{K}_{t}(x).$ We wish to bound $R(T)=(1/T)\sum_{t=1}^{T}\left(\bar{f}_{t}(x^{K}_{t})-\bar{f}_{t}(x_{t})\right),$ which can be rewritten as
$R(T,K)$ $=\frac{1}{T}\sum_{t=1}^{T}\left(\bar{f}_{t}(x^{K}_{t})-\bar{f}_{t-1}(x_{t-1})\right)+\frac{1}{T}\sum_{t=1}^{T}\left(\bar{f}_{t-1}(x_{t})-\bar{f}_{t}(x_{t})\right).$
Note that the second summation telescopes to $(1/T)\left(\bar{f}_{0}(x_{0})-\bar{f}_{T}(x_{T})\right).$ We then begin by bounding the terms in the first summation. We observe that
$\bar{f}_{t}(x^{K}_{t})\leq\bar{f}^{K}_{t+1}(x^{K}_{t})+\min\left\{\Phi^{K}_{t},\max_{j\leq J}M_{j}(2\varepsilon_{t}+d^{K}_{t,1})\right\}=\bar{f}^{K}_{t+1}(x^{K}_{t})+\underline{\psi}_{t+1}^{K},$
using [56, Theorem 5] and Kantorovich-Rubinstein duality. Using the same theorems, we observe $\bar{f}_{t-1}(x_{t-1})\geq\bar{f}_{t}^{K}(x_{t-1})-\bar{\psi}_{t}^{K},$ and by the optimality of $x^{K}_{t}$ for the particular problem, $\bar{f}_{t}^{K}(x_{t-1})\geq\bar{f}_{t}^{K}(x^{K}_{t})$. Furthermore, by Kantorovich-Rubinstein duality
$\bar{f}^{K}_{t+1}(x^{K}_{t})-\bar{f}_{t}^{K}(x^{K}_{t})\leq\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1})).$
Combining these relations, we obtain
$\frac{1}{T}$ $\sum_{t=1}^{T}(\bar{f}_{t}(x^{K}_{t})-\bar{f}_{t-1}(x_{t-1}))\leq\frac{1}{T}\sum_{t=1}^{T}(\underline{\psi}_{t+1}^{K}+\bar{\psi}_{t}^{K}+\max_{j\leq J}M_{j}(2\varepsilon_{t-1}+W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1})).$ (18)
For the final term, we show again by Kantorovich-Rubinstein duality
$\bar{f}_{0}(x_{0})-\bar{f}_{T}(x_{T})\leq\bar{f}_{0}(x_{T})-\bar{f}_{T}(x_{T})\leq\max_{j\leq J}M_{j}(\varepsilon_{0}+\varepsilon_{T}+W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})).$
We use this to prove Theorem 4.1 and its corollaries.
Proof of Theorem 4.1 and Corollary 4.1.1.
We examine the terms in $R(T,K)$ individually, for a finite value $T\gg\tau$. As we assume the Lipschitz condition, we use the corresponding terms in $\underline{\psi}^{K}_{t-1}$ and $\bar{\psi}^{K}_{t}$ as the upper bounds. By assumption, $\varepsilon_{t}\sim(\log(\beta_{t}^{-1})/t)^{1/d}$, so their averages are $O((\log(\beta_{T}^{-1})/T)^{1/d}).$ Note that by the summability of $\beta_{t}$, the convergence rate of $\beta_{t}$ is at least sublinear. Next, recall by the triangle inequality,
$d^{K}_{t,1}\leq W_{1}(\hat{{\mathbf{P}}}_{t},{\mathbf{P}})+W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})+W_{1}({\mathbf{P}}^{K}_{\star},{\mathbf{P}}^{K}_{t}).$
By Theorem 3.1, we can construct a summable sequence $\beta_{t}$ such that ${\mathbf{P}}(W_{1}(\hat{{\mathbf{P}}}_{t},{\mathbf{P}})\leq\varepsilon_{t})\geq 1-\beta_{t}$, where $\varepsilon_{t}$ is the same as above. By Theorem 3.4, a similar result holds for $W_{1}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t})$, with a sequence of radii $\tilde{\varepsilon}_{t}\leq O(\varepsilon_{t})$ for $t\geq\tau$. Since $T\gg\tau$, the finite values $\tilde{\varepsilon}_{t}/T$ for $t\leq\tau$ are dominated by $O(\varepsilon_{T})$. Therefore, the terms involving $d^{K}_{t,1}$ and $d^{K}_{t-1,1}$ are together $O(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star})+\left(\log(\beta_{T}^{-1})/T\right)^{1/d})$, with probability at least $\Pi_{t=0}^{T-1}(1-\beta_{t})^{2}$. This probability can be constructed to converge to a nonzero value due to the summability of $\beta_{t}$. Next, we note again by the triangle inequality,
$W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})\leq W_{1}(\hat{{\mathbf{P}}}^{0},{\mathbf{P}})+W_{1}({\mathbf{P}},\hat{{\mathbf{P}}}_{t}),$
where the first term is constant in $T$ and the second term, by measure concentration results, is upper bounded by $\varepsilon_{T}$ with some probability $1-\beta_{T}$. Therefore, with probability $1-\beta_{T}$, $W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})/T$ is $O(T^{-1})$. Lastly, by the triangle inequality,
$W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1})\leq W_{1}(\hat{{\mathbf{P}}}^{K}_{t},{\mathbf{P}}^{K}_{\star})+W_{1}({\mathbf{P}}^{K}_{\star},\hat{{\mathbf{P}}}^{K}_{t-1})\leq 2\tilde{\varepsilon}_{t-1},$
with probability $(1-\beta_{t-1})^{2}$. The average over $T$ is again $O(\varepsilon_{T})$, with probability $\Pi_{t=0}^{T-1}(1-\beta_{t})^{2}$. Combining all terms, we have $R(T,K)\leq O(\left(\log(\beta_{T}^{-1})/T\right)^{1/d})+O(W_{1}({\mathbf{P}},{\mathbf{P}}^{K}_{\star}))$, with a consolidated probability $1-\zeta\geq\Pi_{t=0}^{T}(1-\beta_{t})^{4}$. Again, we remark that the probability $1-\zeta$ can be constructed to be high, with well-chosen sequences $\beta_{t}$ and $\varepsilon_{t}$. Note that their convergence rates are inversely related; the higher the desired probability $1-\zeta$, the slower the convergence with respect to $T$. Using the Bonferroni Inequality [12], the probability can be estimated as $1-\zeta\geq 1-\sum_{t=0}^{T}4\beta_{t}$.
Corollary 4.1.1 follows by setting $\beta_{t}=O(\exp(-\sqrt{n_{t}})$ and simplifying. ∎
Proof of Corollary 4.1.2.
The asymptotic result for $T\rightarrow\infty$ follows from Lemma 4.1 and the results from Section 3. As $t\to\infty$ and $T\to\infty$, by the almost sure convergence of $\underline{\psi}_{t}^{K}+\bar{\psi}_{t}^{K}$ to $\underline{\psi}_{\star}^{K}+\bar{\psi}_{\star}^{K}$, $W_{1}(\hat{{\mathbf{P}}}^{0},\hat{{\mathbf{P}}}_{t})/T$ to $W_{1}(\hat{{\mathbf{P}}}^{0},{{\mathbf{P}}})/T$ to 00, and $W_{1}(\hat{{\mathbf{P}}}^{K}_{t},\hat{{\mathbf{P}}}^{K}_{t-1})$ to 00, by the Cesàro Mean Theorem [15], $R(T,K)$ converges to some value that is $O(\underline{\psi}_{\star}^{K}+\bar{\psi}_{\star}^{K})$ with probability 1. The result for $K\rightarrow\infty$ follows from Theorem 3.2. ∎
A.3 Proofs of Section 5 results
Proof of Lemma 5.1.
Let the center of the $k$-th cluster be $a^{k}$. By definition of the Wasserstein metric,
$d_{W_{p}}(K,t)^{p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{\mathbf{P}}^{K}_{t})^{p}$ $=\inf_{\pi\in\Pi}\left\{\int_{{S}^{2}}\|u-u^{\prime}\|^{p}\pi(du,du^{\prime})\right\}$
@@ A.3 Proofs of Section 5 results
where the final inequality follows from Assumption 6.1. Next, by exploiting the triangular inequality, with probability $1-\beta$, we get

### v2

@@ 7 Numerical examples
  • Online clustering (our memory-efficient method). We use our online clustering algorithm, as described in Algorithm 3, to update cluster assignments. This method discards the datapoints once seen, and is thus memory-efficient. Upon constructing the empirical clustered distribution, we solve (9) for the online solution.
  • $k$-means reclustering (our method). Same as above, except using $k$-means clustering at each time step, warm-starting at the previous centers. The clustering procedure is described in Algorithm 2.
  • Wasserstein DRO. At time $t$, we solve (9) with the ambiguity set $\mathcal{P}_{t}$.
  • Sample average approximation SAA. At time $t$, we solve the stochastic optimization problem with respect to the empirical measure $\hat{{\mathbf{P}}}_{t}$ of the training data. We denote this in-sample value as $H^{\rm SAA}_{t}$.
We use a training dataset of up to $T=2000$ time steps (datapoints) for all approaches. For evaluation, we employ both validation and testing datasets, comprising of $N^{\rm val}=N^{\rm test}=200$ datapoints. To limit the total solving time, particularly for the nominal DRO problem, we solve the problems every 100 time steps. We compare the following metrics, with values averaged over 30 repetitions of each experiment. We additionally show the 25-75th percentiles using shaded regions.
  • In-sample certificate. We compute the in-sample certificates obtained with respect to the training data. For nominal Wasserstein DRO, we use the in-sample objective $H_{t}$. For our online methods, we use $H^{K}_{t}+\underline{\psi}_{t}^{K}$ as defined in Theorem 3.6. For all DRO approaches, we do not include the theoretical residual $\rho_{t}$, as the radii $\varepsilon_{t}$ are empirically determined as outlined in Section 6.
  • Out-of-sample objectives. For all methods, we also compute the out-of-sample expected value, which is the stochastic optimization objective computed using the empirical measure of the testing dataset, at the given solutions.
  • Empirical confidence. The empirical confidence, denoted $1-\hat{\beta}_{t}$, is computed as the probability that the certificate upper bounds the out-of-sample value.
@@ 7.1 Sparse portfolio optimization
We consider a market that forbids short-selling and has $d$ assets [54, 33]. The uncertain parameters are the daily returns, given by $v\in{\bf R}^{d}$. The decisions $x\in{\bf R}^{n}$} represent the percentage weights (of the total capital) invested in the assets. We restrict our selection to at most $\gamma$ assets, given by the 0-th norm cardinality constraint. The distribution $\mathbf{P}$} is unknown, but we have access to a streaming dataset $\mathcal{D}_{t}$, updated each day with a new returns vector. Our objective is to minimize the $\mathop{\bf CVaR}$ with respect to $x$,
$\begin{array}[]{ll}\text{minimize }&\mathop{\bf CVaR}(-v^{T}x,\omega)\\ \text{subject to}&\mathbf{1}^{T}x=1,\quad x\geq 0,\quad\|x\|_{0}\leq\gamma,\end{array}$
which represents the average of the $\omega$ largest portfolio losses that occur. In other words, the $\mathop{\bf CVaR}$ term seeks to ensure that the expected magnitude of portfolio losses, when they occur, is low. The objective can be written as the expectation of the maximum-of-affine functions [51], i.e., ${\mbox{\bf E}}_{{\mathbf{P}}}\left[\tau+(1/\omega)\max\{-v^{T}x-\tau,0\}\right]$. We solve the online DRO problem using the reformulation (9), with $r=1$ and $\|\cdot\|$ the $\ell_{2}$-norm. As $f$ is maximum-of-affine, $\bar{\psi}_{t}^{K}=0$, and $\Phi^{K}_{t}$ converges to $\Phi^{K}_{\star}$, as proven in Theorem 3.9.
@@
We take stock data from the past 5 years of S&P500 (1/1/2020 to 1/1/2025) daily returns, and generate synthetic data from their fitted general Pareto distributions, with correlations preserved using a Gaussian copula. These distributions are chosen to model the data more accurately than a Gaussian fit. We let $\omega=20\%$, $d=50$ stocks, and restrict our portfolio to at most $\gamma=8$ stocks. We initialize with a dataset of size $n_{0}=5$, and show results for $K=15$ and $K=25$. This range is obtained using the elbow method on the initial dataset, and allowing for adjustments.
@@ Results
In Figure 3, we show the various certificates given in equation (17) (disregarding $\rho_{t}$ in favor of cross-validation), and note that the relationships follow the hierarchy presented. The data-compressed in-sample objective value for this maximum-of-affine problem is lower than the in-sample objective of nominal DRO, but adding the clustering discrepancy makes it a valid certificate. We also show the clustering values for a longer horizon, up to $T=10000$, with $\tau=8500$. We observe that the $p=1$ and $p=2$ clustering distances, $W^{K}_{t,1}$ and $(D^{K}_{t,2})^{2}$, as well as $\Phi^{K}_{t}$, are converging. Moreover, $W^{K}_{t,1}\geq\Phi^{K}_{t}$ and, hence, $\max_{j\leq J}M_{j}W^{K}_{\star,1}\geq\Phi^{K}_{t}$ where the maximum Lipschitz constant is $M_{1}=1/\omega=5$. This illustrates the relative optimality of $\Phi^{K}_{t}$, as opposed to the Wasserstein-1 distance, as proven in Theorem 3.10. In fact, although we have not assumed (or achieved) optimal clustering-induced coupling, the result still held.
Figure 3: Sparse portfolio, $K=25$. The radius $\varepsilon_{t}=0.0025{(t+n_{0})^{-1/40}}$ for all methods. Left: certificates from (17). Right: clustering distances given in Definition 3.4, with $T=10000$. The lighter (red) shade is for the reclustering method.
@@ 8 Conclusions
We have introduced an online data compression framework for solving Wasserstein DRO problems with streaming data. Our method constructs adaptive ambiguity sets using online clustering, allowing the uncertainty model to evolve as new data arrives, while maintaining out-of-sample performance guarantees. We analyzed the impact of data compression on solution quality, providing finite-sample and asymptotic performance bounds that relate to nominal bounds for Wasserstein DRO. The framework is compatible with a broad class of clustering algorithms and supports efficient, memory-aware implementations. Numerical experiments in sparse portfolio optimization demonstrate significant computational savings with minimal loss in solution quality.
@@ Acknowledgments
I. Wang and B. Stellato are supported by the NSF CAREER Award ECCS-2239771. I. Wang is also supported by the Princeton Wallace Memorial Fellowship. B. Stellato is also supported by the ONR YIP Award N000142512147. M. Fochesato is supported by G-Research. The authors are pleased to acknowledge that the work reported on in this paper was substantially performed using the Princeton Research Computing resources at Princeton University.
@@ Appendix A Appendices
For all $t$, let $\tilde{\mathbf{Q}}_{t}\in\tilde{\mathbf{B}}_{\varepsilon_{t}}^{r}(\hat{{\mathbf{P}}}_{t})$ and $\mathbf{Q}_{t}\in\mathcal{P}_{t}$. By [38, Lemma 3.7], we observe ${\mathbf{P}}^{\infty}\{\lim_{t\to\infty}\tilde{\mathbf{Q}}_{t}\rightarrow{\mathbf{P}}\}=1$ and ${\mathbf{P}}^{\infty}\left\{\lim_{t\to\infty}\mathbf{Q}_{t}\rightarrow{\mathbf{P}}\right\}=1$. The first result follows. If $f$ is Lipschitz, by the Kantorovich-Rubinstein duality [30], and by noting $W_{1}(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})\leq W_{r}(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})\leq 2\varepsilon_{t}$ (by the ordering of Wasserstein distances and the triangle inequality), we obtain $\Delta_{t}\leq 2\max_{j\leq J}M_{j}\varepsilon_{t}$. If $f$ is $L$-smooth, then by the descent lemma and (3), $\mathbf{E}_{\mathbf{\tilde{Q}_{t}}}[f(u,x_{t})]-\mathbf{E}_{\mathbf{{Q}_{t}}}[f(u,x_{t})]\leq\inf_{\pi\in\Pi(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})}\int_{S\times S}\nabla f(v,x_{t})^{T}(u-v)d\pi(u,v)+(L/2)W_{2}(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})^{2}.$ By the Cauchy-Schwartz inequality and Hölder’s inequality with order 2, the integral is upper bounded by $(\textstyle\int\|\nabla f(v,x_{t})\|^{2}d{\mathbf{Q}}_{t})^{1/2}(\int\|u-v\|^{2}d\pi(u,v))^{1/2}=\|\nabla f_{x_{t}}\|_{L_{2}({\mathbf{Q}}_{t})}W_{2}(\tilde{\mathbf{Q}}_{t},\mathbf{Q}_{t})$. When $r\geq 2$, this simplifies to $\Delta_{t}\leq\|\nabla f_{x_{t}}\|_{L_{2}({\mathbf{Q}}_{t})}2\varepsilon_{t}+2L\varepsilon_{t}^{2}$. The result follows.
@@
We first prove (15). We begin by observing
$H_{t}=\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\Phi^{K}_{t},$
with $\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]=H^{K}_{t}$ in the final clause. The first inequality follows from the optimality of $x_{t}$, the second follows from [54, Theorem 5], and $\Phi^{K}_{t}$ is defined in Definition 3.4. From [54, Theorem 4], we note that $\Phi^{K}_{t}=0$ for concave functions ffi.e., when $J=1$. When $f$ is Lipschitz, using Kantorovich-Rubinstein duality [30], we can obtain a second bound on $H_{t}$,
$\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]$ $\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\max_{j\leq J}M_{j}(2\varepsilon_{t}+W_{1}(\hat{{\mathbf{P}}}_{t},\hat{{\mathbf{P}}}^{K}_{t})).$
When $f$ is $L$-smooth, using the same procedure as in the proof of Lemma 3.2 above, we obtain a third bound in terms of the $W_{2}$ distance. Combining all gives the bound on $H_{t}-H^{K}_{t}$. For the other side, from the optimality of $H^{K}_{t}$ and [54, Theorem 5],
$H^{K}_{t}\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]$ $\leq H_{t}+(\tilde{H}_{t}-H_{t})+{\max_{j\leq J}\frac{L_{j}}{2}}(D^{K}_{t,2})^{2},$
where $\tilde{H}_{t}-H_{t}=\Delta_{t}$ is given in Lemma 3.2. Kantorovich-Rubinstein duality yields
$H^{K}_{t}\leq\max_{\mathbf{Q}\in\mathcal{P}^{K}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x_{t})]\leq H_{t}+\max_{j\leq J}M_{j}(2\varepsilon_{t}+W^{K}_{t,1}).$
Combining everything, we obtain (15). Using these relationships, and the finite-sample performance guarantee (5) implied by the chosen radii sequence in Assumption 3.2(1), we obtain (17). To prove (16), we note that if the radii are chosen using Theorem 2.1, then with probability $1-\beta_{t}$, ${\mathbf{P}}\in\mathcal{P}_{t},$ and we have, with $\rho_{t}=0$,
${\mathbf{P}}^{{t}}\left(H_{\star}\leq\mathbf{E}_{\mathbf{P}}[f(u,x^{K}_{t})]\leq\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]+\rho_{t}\right)\geq 1-\beta_{t}.$ (20)
Applying the bounds we derived for $\max_{\mathbf{Q}\in\mathcal{P}_{t}}\mathbf{E}_{\mathbf{Q}}[f(u,x^{K}_{t})]$ gives the desired result. If the radii are chosen using Theorem 2.2, then as $x^{K}_{t}\in\mathcal{X}$, (20) holds by (6).
A.2 Proofs of Section 4 results
@@
By definition of the Wasserstein metric,
$(W^{K}_{t,p})^{p}=W_{p}(\hat{{\mathbf{P}}}_{t},\hat{\mathbf{P}}^{K}_{t})^{p}$ $=\inf_{\pi\in\Pi}\left\{\int_{{S}^{2}}\|u-u^{\prime}\|^{p}\pi(du,du^{\prime})\right\}$
@@
where the final inequality follows from the cluster diameter of $2\eta_{K}$, using Definition 3.5. Next, by exploiting the triangular inequality, with probability $1-\beta$, we get